OpenStudy (freemap):

Which ordered pairs are solutions to the inequality x + 3y ≥ –8? Choose all answers that are correct. A.(–5, –1) B.(0, –3) C.(–1, –2) D.(–6, 0) E.(–16, 2)

OpenStudy (texaschic101):

this is just a matter of subbing and seeing which ones are true. Lets try the first one... x + 3y >= -8 (-5,-1)....x = -5 and y = -1 now we sub -5 + 3(-1) >= -8 -5 - 3 >= -8 -8 >= -8 --- correct because -8 = -8 so this is a solution. You want to try the next one ?

OpenStudy (freemap):

YES

OpenStudy (texaschic101):

x + 3y >= -8 (0,-3)...x = 0 and y = -3 now we sub 0 + 3(-3) >= -8 can you finish and see if it is true ?

OpenStudy (freemap):

-9 is less than -8 so false

OpenStudy (texaschic101):

wait...your sign is greater then or equal It is saying -9 is greater then or equal to -8....so it is false

OpenStudy (freemap):

sorry just looked at sign so it'll still be false because -9 is not greater than -8

OpenStudy (texaschic101):

keep this in mind....with positive numbers, the bigger the number, the greater it is...but with negatives, the smaller number is greater

OpenStudy (freemap):

Understood

OpenStudy (texaschic101):

want to try the next one because the 2nd one is not a solution

OpenStudy (freemap):

yes I'll try, just a sec

OpenStudy (freemap):

5>-8 which is true

OpenStudy (texaschic101):

x + 3y >= -8...x = -1 and y = -2 -1 + 3(-2) >= -8 -1 - 6 >= -8 -7 >= -8......this is true

OpenStudy (freemap):

I was just about to submit -7>=-8 which is true. Instead of me multiplying 3 *-2 I did 3*2

OpenStudy (texaschic101):

lol...be careful of your signs....they can mess with you every time

OpenStudy (freemap):

Every single time lol, about to work out the next one

OpenStudy (texaschic101):

ok...you know how to do this..you have proven that...now let see what you can do :)

OpenStudy (freemap):

-6>-8 is true

OpenStudy (texaschic101):

correct ^^ I told you that you could do this...great job

OpenStudy (freemap):

Thanks lol, Last one here I go

OpenStudy (freemap):

-10>-8 is false

OpenStudy (texaschic101):

you did it again...correct....woot woot...doing the happy dance now :)

OpenStudy (freemap):

lol Me to Celebration time, Thanks so much means so much that you'd stay even though I'm slow. I wish I could be like you just work problems out in a second without help. Thankssss

OpenStudy (texaschic101):

it takes practice and patience. You will one day be helping others learn this stuff. So your solutions are A,C,D because B and E were false....but you already knew this so why am I telling you...lol

OpenStudy (freemap):

lol I'm so happy. Thank you so much. I have one more question I think its A, its a graph though uggh! would you help, I'll post again?

OpenStudy (texaschic101):

you can post on this post....but I can't make any promises about knowing the answer....graphs are not my friend...lol

OpenStudy (freemap):

Thanks for helping

OpenStudy (freemap):

Which is the graph of y ≤ 1 – 3x?

OpenStudy (freemap):

I said A because it has a solid line, and its less than so it goes to the left

OpenStudy (texaschic101):

y = -3x + 1 --- I just rearranged it and temporarily removed the < sign y intercept is 1 or (0,1) x intercept can be found by subbing in 0 for y.. 0 = -3x + 1 3x = 1 x = 1/3 ... or (1/3,0) now back to your signs....it contains an equal sign, so it is a solid line. It is < (less then) so shading occurs below (or to the left) of the line. the correct graph is : A...you are correct :)

OpenStudy (texaschic101):

it takes me a little longer with graphs...lol

OpenStudy (freemap):

Never a problem, Thanks you're a hero. I'm going to take what you taught and apply to my other ones. Have a great night. I always like getting help from you, so great at explaining complicated things. Thanks so so much

OpenStudy (texaschic101):

glad to help :) good luck

OpenStudy (freemap):

Thank you

OpenStudy (texaschic101):

:)